OpenStudy (anonymous):

(8x^4+6x^3+2x^2) divided by (2x^2)

OpenStudy (whpalmer4):

Divide each term individually, then add the results.

sam (.sam.):

Split and divide, \[\frac{8x^4+6x^3+2x^2}{2x^2}=\frac{8x^4}{2x^2}+\frac{6x^3}{2x^2}+\frac{2x^2}{2x^2}\]

OpenStudy (whpalmer4):

Yes, if you do what we told you, you'll have the answer.

OpenStudy (whpalmer4):

OpenStudy is not intended as an answer source.

OpenStudy (whpalmer4):

Happy to check your work, however.

OpenStudy (anonymous):

4x^2+3x+1

OpenStudy (whpalmer4):

If you just copy .Sam.'s answer, you probably aren't going to get it marked correct, he hasn't completed the problem, just written out what I told you.

OpenStudy (anonymous):

No. it's an online class so, I got it correct